OpenStudy (anonymous):

The altitude of a triangle is 5 m less than its base. The area of the triangle is 42 m^2. Find the length of the base.

OpenStudy (anonymous):

LOL very nice h = b - 5 (1/2)*b*h = area (1/2)*b*h = 42 b(b - 5) = 84 b^2 - 5b - 84 = 0 (b + 7)(b - 12) b = -7, 12 b has to be positive because it's the length so be is 12

OpenStudy (anonymous):

oh okay i get now thank you

OpenStudy (anonymous):

no problem